# DreamMix: Decoupling Object Attributes for Enhanced Editability in Customized Image Inpainting

### v1

@@ 3.4 Textual Attribute Substitution
The attribute decoupling mechanism employs advanced VLMs to enhance the diversity of training samples, thereby effectively adapting pretrained inpainting models for subject-driven inpainting. However, due to the lack of unseen attribute words during training, relying solely on attribute decoupling mechanism still poses challenges, especially when the target attributes differ significantly from the object identity. To address this, we introduce a Textual Attribute Substitution (TAS) module during the testing phase to further mitigate the influence of object identity for more precise attribute editing. The pipeline of TAS is illustrated in Fig. 2 (b).
@@
$\boldsymbol{p}_{dec}=\boldsymbol{p}_{raw}-\left(\frac{\boldsymbol{p}_{raw}\cdot\boldsymbol{p}_{eli}}{\|\mathbf{p}_{eli}\|}\right)\frac{\boldsymbol{p}_{eli}}{\|\boldsymbol{p}_{eli}\|},$ (9)
where $\boldsymbol{p}_{raw}\cdot\boldsymbol{p}_{eli}$ denotes the dot product of $\boldsymbol{p}_{raw}$ and $\boldsymbol{p}_{eli}$, and $\|\boldsymbol{p}_{eli}\|$ is the norm of $\boldsymbol{p}_{eli}$. After applying this embedding substitution, the conflicting features of the original object identity are effectively decoupled, making the inpainting model focus on the demand of the target prompt. As shown in Fig. 2 (c), we apply the decomposed text embedding $\boldsymbol{p}_{dec}$ to the cross-attention layers at both LCG and GCH stages to boost the editing efficacy of the target text.
@@ 4 Experiments
Method Identity Preservation Attribute Editing
CLIP-T CLIP-I DINO FID CLIP-T CLIP-I DINO FID
Large-scale Training Paint-by-Example [39] 0.238 0.541 0.582 44.6 - - - -
MimicBrush [5] 0.256 0.614 0.631 47.0 - - - -
AnyDoor [6] 0.284 0.688 0.711 53.7 - - - -
IP-Adapter [40] 0.268 0.633 0.633 45.3 0.236 0.611 0.629 47.5
LAR-Gen [26] 0.269 0.662 0.676 48.7 0.257 0.656 0.668 49.6
Tuning-free TIGIC [18] 0.262 0.533 0.584 45.2 0.236 0.479 0.548 47.3
Few-shot Fine-tuning DreamBooth 0.248 0.609 0.604 40.8 0.235 0.567 0.595 42.0
DreamBooth [33] 0.253 0.644 0.628 40.5 0.235 0.601 0.612 41.7
DreamMix 0.284 0.713 0.685 43.4 0.284 0.659 0.672 44.2
DreamMix 0.285 0.728 0.712 42.9 0.289 0.674 0.695 43.9
Real Image 0.315 0.883 0.865 35.9 - - - -
Table 1: Quantitative comparison of different methods on Identity Preservation and Attribute Editing. The “-” symbol indicates that the method don’t support text-driven editing. We divide the compared methods into three types based on their training protocol: large-scale training, tuning-free, and few-shot finetuning. $\dagger$” indicates that only one image is used in model finetuning for each subject.
Figure 4: Visual comparison between different methods. From left to right are input image and subject, visual results of our DreamMix, IP-Adapter [40], DreamBooth [33], TIGIC [18], AnyDoor [6], and LAR-Gen [26].
@@ 4.1 Experiment Setup
Implementation Details. We use Fooocus inpainting model [20], a variant of SDXL [28], as our base generator. A LoRA [15] with a rank of 4 is applied to $W_{k}$ and $W_{v}$ in the attention layers. During finetuning, images are resized to $1024\times 1024$, and we use the Adam optimizer with an initial learning rate of $41e-$ 4. The total finetuning time for each subject is proximately 20 minutes on a single RTX 4090 GPU. The hyperparameter $\lambda$ in DIF is set to 0.7.
@@
Benchmarks. We follow prior works [6, 26, 18] to use 30 subjects from DreamBooth [33] for performance evaluation. We construct a benchmark that contains over 4,000 high-quality image-mask pairs. Specifically, we collect the images from the COCO-val2017 dataset [19], filtering out images or bounding boxes with lower resolutions. We mix the 30 subjects with these background images, resulting in an average of approximately 140 background images per subject. Besides, we create 3–5 prompts for identity preservation or attribute editing for each object.
@@
Evaluation Metrics. Our method employs both image and text guidance for simultaneously object composting and text-driven editing. To evaluate subject identity similarity, we follow DreamBooth [33] to compute CLIP-I [29] and DINO [25] metrics between the generated images and the source images. To assess the efficacy of attribute editing, we compute the CLIP-T metric to measure the cosine similarity between the generated images and the text prompts. To evaluate the realism of generated images, we measure the Fréchet Inception Distance (FID) [12] using the COCO-test2017 dataset as real image distribution reference. Additionally, we conduct user studies as a subjective evaluation metric for our method.
4.2 Comparison with Other Methods
@@
To highlight the efficacy of our method on identity-preserved object compositing and text-driven attribute editing, we conduct comparison experiments on these two tasks separately. The quantitative and qualitative comparison results are illustrated in Tab. 1 and Fig. 4. We select a range of classic and recent methods and categorize them into three types: large-scale training, tuning-free, and few-shot finetuning. For most of the compared methods [6, 5, 39, 26, 18], we exploit their open-source models to obtain visual results. Additionally, we apply IP-Adapter [40] and DreamBooth [33] on our base inpainting model [20] to better verify our method’s efficacy.
Identity Preservation. For identity preservation, we evaluate the “Real Image” as an upper bound for all methods, where CLIP-T, CLIP-I and DINO are calculated on the training dataset, and FID is measured on our benchmark images. As shown in Tab. 1, the proposed DreamMix outperforms both large-scale training and tuning-free approaches across all metrics. Compared to DreamBooth, our method surpasses it on three metrics while is slightly lower in FID. The possible reason is that DreamBooth tends to produce ID-irrelevant fill-in results that remain harmonious within the scene (see the fourth row in Fig. 4). Additionally, the visual comparison in the first two rows of Fig. 4 demonstrates that our model is effective at producing identity-preserved, context-harmonious inpainting results.
@@
Attribute Editing. Since some compared methods [6, 5, 39] don’t support text-driven editing, we can’t report their metrics on this task. As shown in Tab. 1, our method outperforms the other compared approaches on CLIP-I, CLIP-T, and DINO, demonstrating that our model can accurately generate the editing effects of text prompts while preserving the original identity. As shown in the last three rows of Fig. 4, our method supports diverse types of attribute editing, such as color, shape, and adding accessories. More visual results are provided in the supplementary materials.
@@
User Study. We conduct a user study to further evaluate the feasibility of our method. For both tasks, we ask 100 volunteers to complete a questionnaire consisting of 30 questions, resulting in a total of 3000 responses. In each question, we present an image of a reference object, a background image, a target text, and the generated images of anonymous methods. Participants are asked to answer the following question: “Which of the following images is the best considering image quality and similarity to the subject, and match with text description?” We summarize the results and present them in Tab. 2. As we can see, the proposed DreamMix has an overwhelming preference for both identity preservation and attribute editing.
@@ 4.3 Ablation Studies
We conduct ablation studies on both identity preservation and attribute editing tasks to verify the efficacy of Disentangled Inpainting Framework (DIF), Textual Attribute Substitution (TAS) module, and Attribute Decoupling Mechanism (ADM). The quantitative results are presented in Tab. 3, and the qualitative results are shown in Fig. 5 and Fig. 6.
Method Identity Preservation Attribute Editing
IP-Adapter [40] 8% 7%
MimicBrush [5] 12% -
AnyDoor [6] 10% -
TIGIC [18] 2% 4%
LAR-Gen [26] 13% 15%
DreamMix 55% 74%
Table 2: User study results for Identity Preservation and Attribute Editing tasks, showing user preference percentages for each method. DreamMix outperforms other methods significantly in both Identity Preservation (55%) and Attribute Editing (74%), indicating strong user preference.
@@
Identity Preservation Baseline 0.253 0.644 0.628
+DIF 0.291 0.723 0.717
++TAS 0.290 0.728 0.715
$+$ADM 0.285 0.728 0.712
Attribute Editing Baseline 0.235 0.601 0.612
@@
$+$ADM 0.289 0.674 0.695
Table 3: Quantitative comparison of our ablation methods. The Baseline method indicates the base inpainting model [20] combined with DreamBooth’s finetuning strategy [33].
@@
Effect of $\lambda$ in DIF. In DIF, we use $\lambda$ to control the separation of local content generation and global context harmonization. We conduct experiments with varying $\lambda$ values in our model and present the visual results in Fig. 5. As we can see, as $\lambda$ increases, the inpainted regions become more harmonious with the context, but the allocation accuracy may decrease. Using only GCH (i.e., $\lambda=1$) may pose generation failures (e.g., clock case) when inpainting on small sizes. Conversely, relying solely on local inpainting (i.e., $\lambda=0$) can affect the overall harmony of the inpainted regions. We find that $\lambda=0.7$ provides an optimal balance in most cases.
Effectiveness of Each Component. We conduct experiments to verify the effectiveness of the three core components in our DreamMix. We use DreamBooth [33] on our base inpainting model [20] as a baseline method. Then, we incrementally incorporate each component and evaluate them on both tasks. The results are illustrated in Tab. 3. As we can see, the proposed DIF effectively improves the generative capacity of the baseline method in both tasks. For attribute editing, our model achieves a continual increase in CLIP-T as the incremental involve of TAS and ADM, while exhibiting a slight decrease in CLIP-I and DINO. These results indicate that the edited regions are more consistent with the text prompts and, consequently, slightly differ from the original appearance in the feature space. This demonstrates the effectiveness of the proposed TAS and ADM in improving text-driven editing. In addition, the results in identity preservation demonstrate that incorporating TAS and ADM does not affect identity preservation if no editing prompts are given. More importantly, all our intermediate results surpass the compared methods in Tab. 1, further demonstrating the efficacy of the proposed modules. Fig. 6 illustrates the visual examples of ablation experiments. In the first row, we observe that DIF enhances image quality, while TAS and ADM do not negatively impact the subject’s identity. In the second row, DIF establishes a solid foundation that allows TAS and ADM to effectively improve editing precision.

### v2

@@ 3.3 Textual Attribute Substitution
The attribute decoupling mechanism employs advanced VLMs to enhance the diversity of training samples, thereby effectively adapting pretrained inpainting models for subject-driven inpainting. However, due to the lack of unseen attribute words during training, relying solely on attribute decoupling mechanism still poses challenges, especially when the target attributes differ significantly from the object identity. To address this, we introduce a Textual Attribute Substitution (TAS) module during the testing phase to further mitigate the influence of object identity for more precise attribute editing. The pipeline of TAS is illustrated in Figure 3 (b).
@@
$\boldsymbol{p}_{dec}=\boldsymbol{p}_{raw}-\left(\frac{\boldsymbol{p}_{raw}\cdot\boldsymbol{p}_{eli}}{\|\boldsymbol{p}_{eli}\|}\right)\frac{\boldsymbol{p}_{eli}}{\|\boldsymbol{p}_{eli}\|},$ (5)
where $\boldsymbol{p}_{raw}\cdot\boldsymbol{p}_{eli}$ denotes the dot product of $\boldsymbol{p}_{raw}$ and $\boldsymbol{p}_{eli}$, and $\|\boldsymbol{p}_{eli}\|$ is the norm of $\boldsymbol{p}_{eli}$. After applying this embedding substitution, the conflicting features of the original object identity are effectively decoupled, making the inpainting model focus on the demand of the target prompt. As shown in Figure 3 (c), we apply the decomposed text embedding $\boldsymbol{p}_{dec}$ to the cross-attention layers at both LCG and GCH stages to boost the editing efficacy of the target text.
3.4 Disentangled Inpainting Framework
In addition to ADM and TAS, we propose a Disentangled Inpainting Framework (DIF) to enhance object insertion in confined local regions. As shown in Figure 3 (c), the DIF consists of two seamless stages of Local Content Generation (LCG) and Global Context Harmonization (GCH), to enhance local subject integration and global visual coherence.
Local Content Generation. To improve the success of small object insertion, we implement LCG at early time steps of $\lambda{T}$, where $\lambda\in[0,1]$ is a trade-off for time step separation. To improve the generation ability across varying-sized areas, we crop the background image $\boldsymbol{x}$ into a local patch and obtain the latent code as $\boldsymbol{z}^{L}=\text{Enc}(\text{Crop}(\boldsymbol{x}_{s},\boldsymbol{m}^{\prime}))$. Here, $\text{Crop}(\cdot)$ denotes an image cropping operation guided by a mask $\boldsymbol{m}^{\prime}$, which is a slightly enlarged region of original inpainting area, similar to the approach used in AnyDoor [5]. $\text{Enc}(\cdot)$ is the VAE encoder. At each time step, the noise predictor $\boldsymbol{\epsilon}_{\theta}$ takes the text prompt $\boldsymbol{p}_{dec}$ as input to generate the denoised latent code $\boldsymbol{z}_{t}$ by $\boldsymbol{z}_{t}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t+1};\boldsymbol{p}_{dec},t)$. During $\lambda{T}$ steps, we exploit a blending strategy [1] to remain the content of non-editing areas. This process can be formulated as:
where $\odot$ denotes element-wise multiplication and $\boldsymbol{z}_{t}^{L}$ is the noisy latent code of $\boldsymbol{x}^{L}$. $\hat{\boldsymbol{z}}_{t}$ denotes the blended latent code that will be fed to the noise predictor $\boldsymbol{\epsilon}_{\theta}$ in the next time step. After the denoising process of $\lambda T$ steps, we project the final result as an RGB image $\hat{\boldsymbol{x}}^{L}$ and repaste it into the original image $\boldsymbol{x}$ to form an intermediate image $\boldsymbol{x}^{G}$, which will be used as an image guidance in the global context harmonization stage.
Global Context Harmonization. The LCG in the preceding time steps effectively yields accurate object compositing with a well-defined layout and object appearance. However, without incorporating the global image content, it tends to produce a disharmonious “copy-paste” effect in the inpainting areas. To address this issue, we implement GCH on the remaining $(1-\lambda)T$ steps to improve the overall harmony between the generated local region and the background image. Similar to LCG, we start by transforming the intermediate image $\boldsymbol{x}^{G}$ into a latent code $\boldsymbol{z}^{G}$ by $\boldsymbol{z}^{G}=\text{Enc}(\text{Crop}(\boldsymbol{x}^{G},\boldsymbol{m}))$. Here, $\boldsymbol{m}$ is the original user-specific mask. Then, we adopt the same blending strategy in GCH as follows:
While DIF uses a blending strategy on explicit input space, its key strength lies in its two-phase optimization mechanism of LCG and GCH. This decoupling prevents the blur-or-conflict trade-off inherent to single-phase blending, ensuring both precise local alignment and seamless global coherence, particularly critical for small-object inpainting.
Figure 5: Visual comparison between different methods. From left to right are input background image and subject, visual results of our DreamMix-SDXL, IP-Adapter-SDXL [61], DreamBooth-SDXL [44], TIGIC [26], AnyDoor [5], and LAR-Gen [37].
@@ 4.1 Experiment Setup
Implementation Details. We apply DreamMix on various inpainting backbones, including SD1.5 [45, 40], SD2.0 [47, 40], SDXL [24], and FLUX [21, 20]. Note that the SDXL backbone used here is an open-source toolkit developed by [24], and we re-implement it using the Diffusers [52] library for fine-tuning. For all backbones, a LoRA [13] with a rank of 4 is applied to $W_{k}$ and $W_{v}$ in the attention layers. The hyperparameter $\lambda$ in DIF is set to 0.7. The VLM used in ADM and TAS is ChatGPT-4o which helps to generate attribute dictionary and formulate 30 text prompts as regularization benchmark.
@@
Benchmarks. We follow prior works [5, 37, 26] to use 30 subjects from DreamBooth [44] for performance evaluation. We construct a benchmark that contains over 4,000 high-quality image-mask pairs. Specifically, we collect images from the COCO-val2017 dataset [25], filtering out images or bounding boxes with lower resolutions. We mix the 30 subjects with these background images, resulting in an average of approximately 140 background images per subject. Besides, we manually create 3–5 prompts for identity preservation or attribute editing for each object. We show some examples of our benchmarks in Figure 6.
@@
Evaluation Metrics. Since our method employs image and text guidance for simultaneous object composting and text-driven editing, we adopt metrics to evaluate image-text correspondence and image-subject similarity. Specifically, to evaluate subject identity similarity, we follow DreamBooth [44] to compute CLIP-I [43] and DINO [33] metrics between generated images and source images. To assess the efficacy of attribute editing, we compute the CLIP-T metric to measure the cosine similarity between generated images and text prompts. Additionally, we conduct user studies as a subjective evaluation metric. To focus on the quality of inpainted regions while considering their integration with the full image, we crop and enlarge the inpainted areas to calculate the evaluation metrics.
4.2 Comparison with State-of-the-art Methods
Table 1: Quantitative comparison of different methods on Identity Preservation and Attribute Editing. The “–” symbol indicates that the method does not support text-driven editing. We label the best and second methods with Bold and Underlined styles, respectively.
Identity Preservation Attribute Editing
Method Backbone Params Dataset CLIP-T CLIP-I DINO CLIP-T CLIP-I DINO
PbE [58] SD1.4 1.3B 1.9M 0.238 0.541 0.593
MimicBrush [4] SD1.5 2.5B 10M 0.256 0.614 0.650
DreamBooth [44] SD1.5 1.3B 4-6 0.251 0.607 0.631 0.239 0.648 0.609
DreamMix SD1.5 1.3B 4-6 0.276 0.693 0.699 0.280 0.653 0.671
TIGIC [26] SD2.1 1.3B 0 0.262 0.533 0.595 0.236 0.479 0.548
AnyDoor [5] SD2.1 2.4B 0.4M 0.284 0.688 0.696
DreamBooth [44] SD2.0 1.3B 4-6 0.258 0.627 0.639 0.240 0.607 0.611
DreamMix SD2.0 1.3B 4-6 0.284 0.697 0.709 0.286 0.668 0.679
IP-Adapter [61] SDXL 5.6B 10M 0.268 0.633 0.649 0.236 0.611 0.629
LAR-Gen [37] SDXL 4.5B 0.6M 0.269 0.662 0.688 0.257 0.656 0.668
DreamBooth [44] SDXL 3.4B 4-6 0.253 0.644 0.641 0.235 0.601 0.612
DreamMix SDXL 3.4B 4-6 0.285 0.728 0.730 0.289 0.674 0.695
ACE++ [30] FLUX 16.8B 0.7B 0.273 0.689 0.712 0.262 0.668 0.674
DreamBooth [44] FLUX 16.8B 4-6 0.261 0.670 0.691 0.253 0.655 0.664
DreamMix FLUX 16.8B 4-6 0.284 0.741 0.728 0.290 0.684 0.701
Real Image 0.315 0.883 0.877
@@ 4.2 Comparison with State-of-the-art Methods
To highlight the efficacy of our method on identity-preserved object compositing and text-driven attribute editing, we conduct comparison experiments on two tasks separately. The quantitative and qualitative comparison results are illustrated in Table 1 and Figure 5. We select a range of classic and recent methods and categorize them based on the backbone versions. For most of the compared methods [5, 4, 58], we exploit their open-source models to obtain visual results. For LAR-Gen [37], as the RefineNet is not publicly available, we evaluate it using only the “Locate” and “Assign” modules on SDXL with an image scale of 1. For TIGIC [26], we follow their official implementation and use SD-2.1 as the base model. For IP-Adapter [61] and DreamBooth [44], we use the same base model as DreamMix to better verify our method’s efficacy. For DreamBooth, we employ the same training strategy as used in DreamMix.
Identity Preservation. For identity preservation, we evaluate the “Real Image” as an upper bound for all methods, where CLIP-T, CLIP-I and DINO are calculated on the training dataset As shown in Table 1, the proposed DreamMix consistently outperforms other methods on different backbones. Anydoor achieves comparable CLIP-T to DreamMix by leveraging large-scale training data and image-guided ID retention to preserve class attributes. However, it struggles with contextual harmonization (lower CLIP-I/DINO) and lacks text-driven editing capabilities, significantly limiting its versatility for dynamic customization tasks. Additionally,the visual comparison in Figure 5 demonstrates that our model outperforms other methods in producing identity-preserved context-harmonious inpainting results.
Figure 6: Examples of our benchmark. Row 1: background images to be edited; Row 2: masks marking the inpainting regions; Row 3: objects to be generated in those regions; Row 4: text prompts describing the subjects.
@@
Attribute Editing. Since some compared methods [5, 4, 58] don’t support text-driven editing, we can’t report their metrics on this task. As shown in Table 1, our method outperforms the other compared approaches on CLIP-I, CLIP-T, and DINO, demonstrating that our model can accurately generate the editing effects of text prompts while preserving the original identity. As shown in the last three rows of Figure 5, our method supports diverse types of attribute editing, such as color, shape, and adding accessories. In contrast, other methods struggle with effective subject editing, leading to compromised identity or leaving the target unchanged. Additionally, DreamMix can use various specific masks to enhance fine-grained pose control over editing regions. More visual results are provided in the supplementary materials.
VLM Metrics in Attribute Editing. We assess attribute-level text–image alignment using the Attribute Binding metric from T2I-CompBench [14], with ChatGPT-4o serving as the vision–language judge. Specifically, for each edited image, the judge is given the text prompt and the edited output and independently rates three dimensions on a 1–5 scale (higher is better): color, shape, and texture consistency. As summarized in Table 2, our method attains strong Attribute Binding scores across all three dimensions, indicating accurate control of color and shape while preserving semantic consistency with the text.
Table 2: Attribute Binding (T2I-CompBench) scores for attribute editing, with ChatGPT-4o judging color, shape, and texture consistency (1–5; higher is better)
Method Backbone Color↑ Shape↑ Textual↑
IP-Adapter SDXL 1.9 2.0 2.0
LAR-Gen SDXL 2.4 2.7 2.5
ACE++ FLUX 2.7 2.9 2.6
DreamBooth SDXL 2.2 2.3 2.3
DreamMix SDXL 3.3 3.6 3.2
@@
User Study. We conduct a user study to further evaluate the feasibility of our method on identity preservation and attribute editing. We ask 100 volunteers to complete a questionnaire consisting of 30 questions, resulting in a total of 3000 responses. In each question, we present an image of a reference object, a background image, a target text, and the images generated by anonymous methods. Participants are asked to answer the following questions: “Which of the following images is the best by considering Image Quality (IQ), Image Similarity (IS) and Text Relevance (TR)?” We summarize the results and present them in Table 3. As shown, the proposed DreamMix has an overwhelming preference for both identity preservation and attribute editing.
Table 3: User study results for Identity Preservation and Attribute Editing tasks, showing user preference percentages for each method. DreamMix outperforms other methods significantly in both categories, indicating strong user preference.
ID Pres. Attr. Editing
Method Backbone IS IQ TR IS IQ
TIGIC SD2.1 3% 2% 1% 2% 1%
LAR-Gen SDXL 19% 21% 13% 12% 13%
IP-Adapter SDXL 4% 8% 5% 6% 8%
DreamBooth SDXL 14% 15% 10% 7% 11%
DreamMix SDXL 60% 54% 71% 73% 67%
@@ 4.3 Ablation Studies
Effectiveness of Each Component. We conduct experiments to verify the effectiveness of Textual Attribute Substitution (TAS) module, Attribute Decoupling Mechanism (ADM) and Disentangled Inpainting Framework (DIF) in DreamMix. The quantitative results are presented in Table 4. We use DreamBooth [44] on our SDXL inpainting model [24] as a baseline method. We incrementally incorporate each component and evaluate them on both tasks. The results in Table 4 demonstrate that the proposed DIF effectively improves the generative capacity of the baseline method in both tasks. For attribute editing, our model achieves a continual increase in CLIP-T as the incremental involve of TAS and ADM, while exhibiting a slight decrease in CLIP-I and DINO. These results indicate that the edited regions are more consistent with the text prompts and, consequently, slightly differ from the original appearance in the feature space. This demonstrates the effectiveness of the proposed TAS and ADM in improving text-driven editing. In addition, the results on identity preservation demonstrate that incorporating TAS and ADM does not affect identity preservation if no editing prompts are given. Figure 7 illustrates the visual examples of ablation experiments. In the first row, we observe that DIF enhances image quality, while TAS and ADM do not negatively impact the subject’s identity. In the second row, DIF establishes a solid foundation that allows TAS and ADM to effectively improve editing precision.
Table 4: Quantitative comparison of our ablation methods. The “Baseline” method indicates the base inpainting model [24] combined with DreamBooth’s finetuning strategy [44]. The “*” symbol indicates single-image training.
@@
ID
Pres.
Baseline 0.253 0.644 0.641
+DIF 0.291 0.723 0.735
++TAS 0.290 0.728 0.732
+++ADM 0.285 0.728 0.730
DreamBooth* 0.248 0.609 0.614
DreamMix* 0.284 0.713 0.711
Attr.
Editing
Baseline 0.235 0.601 0.612
@@
+++ADM 0.289 0.674 0.695
DreamBooth* 0.235 0.567 0.595
DreamMix* 0.284 0.659 0.672
@@
Effect of $\lambda$ in DIF. In DIF, we use $\lambda$ to control the separation of local content generation and global context harmonization. We conduct experiments with varying $\lambda$ values and present the visual results in Figure 8. As we can see, as $\lambda$ increases, the inpainted regions become more harmonious with the context, but the allocation accuracy may decrease. Using only GCH (i.e., $\lambda=1$) may pose generation failures (e.g., clock case) when inpainting on small sizes. Conversely, relying solely on local inpainting (i.e., $\lambda=0$) can affect the overall harmony of the inpainted regions.
Additional Experiments on ADM and TAS Configurations. As shown in Table 5, we conducted a series of additional experiments to evaluate the performance of our system under various ADM and TAS configurations. First, we replaced ChatGPT-4o with alternative VLMs in both ADM and TAS modules. Next, we evaluated the impact of varying the number of regularization images. Finally, we retrained the model using five different randomized attribute orders. Overall, the system showed stable performance across these configuration variations, demonstrating its robustness.
Table 5: Consolidated table showing performance across different VLMs, regularization image counts, and attribute orders.
Category Attri. Edit CLIP-T CLIP-I DINO
VLM Qwen2-VL-2B [53, 39] 0.287 0.675 0.695
LLaVA-v1.5-7B [23] 0.290 0.674 0.692
SmolVLM-0.5B [29, 15] 0.287 0.672 0.694
ChatGPT-4o [34] 0.289 0.674 0.695
Reg. Img. 10 0.285 0.677 0.696
20 0.288 0.676 0.695
30 0.289 0.674 0.695
50 0.290 0.675 0.694
Order random $\Delta$ ±0.002 ±0.004 ±0.002
Table 6: Impact of hyperparameter variations on model performance.
$\tau_{1}$ $\tau_{2}$ $\beta$ CLIP-T CLIP-I DINO
1.5 0.7 0.4 0.289 0.674 0.695
1.2 0.8 0.4 0.288 0.674 0.695
1.0 1.0 0.4 0.286 0.675 0.696
1.5 0.7 0.2 0.288 0.673 0.696
1.5 0.7 1 0.285 0.672 0.693
Hyper-parameter Ablation Studies. The value selected for $\tau_{1}$, $\tau_{2}$ in Eq. 7 and $\beta$ in Eq. 8 are grounded on empirical observations. To further explore their influence, We altering these parameters empirically as detailed in Table 6. Our results demonstrate that while these adjustments had a negligible effect on the final performance metrics, they did impact the model’s convergence speed. Specifically, smaller or larger values of the parameters slightly accelerated or slowed down the convergence, but the overall performance remained consistent across different settings. This suggests that the model is relatively robust to variations in these hyperparameters.
Training on the Single Image. To further validate the effectiveness of DreamMix with fewer training data, we retrain our DreamMix and DreamBooth with just one image per subject. As shown in Table 4, compared to the DreamBooth baseline, DreamMix indicates less performance degradation due to the incorporation of three key components. More importantly, even with only one image, DreamMix remains competitive with state-of-the-art approaches in Table 1.
4.4 Other Applications
Pose Control Using Masks. Fine-grained pose control can be achieved by leveraging specific masks for editing selected regions of an object. Unlike previous inpainting methods [6, 4], which are trained by randomly generating masks, we use SAM to obtain accurate boundary masks of objects and inflate them randomly, which not only greatly accelerates the fitting speed of the model but also gives the model the ability to use the shape of the mask to control the pose of the generated object as shown in Figure 9.
Figure 9: Examples of fine-grained pose control using masks.
Multi-Subject Inpainting. Since our method cannot perform multi-subject inpainting in an end-to-end manner, we composite each subject on an image iteratively to support this task. As shown in Figure 10, this approach allows our method to effectively composite multiple objects into a scene, facilitating object interaction and enabling detailed attribute editing.
Figure 10: Visual examples of multi-subject inpainting, where multiple objects can be seamlessly integrated into a scene with interactive and editable attributes.